# Evolution of the Inner Accretion Flow in Swift J1727.8–1613 across Intermediate States: Insights from Broadband Spectral and Timing Analysis

### v1

@@ Abstract
We present a comprehensive broadband spectral and variability study of the newly detected black hole X-ray binary Swift J1727.8–1613 in the intermediate states during its 2023 outburst, using multi-mission observations from NICER, NuSTAR, AstroSat, and Insight-HXMT. The spectral data up to $78$ keV in the hard-intermediate state (HIMS) requires models with two Comptonizing regions. In contrast, models with a single Comptonizing region adequately describe the soft-intermediate states (SIMS), implying a significant evolution in the disk-corona geometry between the states. The hard X-ray tail above $100$ keV in the HIMS, detected with both AstroSat/CZTI and Insight-HXMT/HE, indicates that the electron population in the corona is not purely thermal but rather hybrid, with a power-law distribution above the thermal cutoff. While both the reflection modeling and disk continuum fitting favor a truncated disk geometry in the HIMS, the disk substantially moves close to the innermost stable circular orbit in the SIMS, accompanied by a significant rise in the disk temperature. This interpretation is further supported by the increase in the QPO frequency from $\sim 1.3$ to $\sim 6.6$ Hz. From joint modeling of the disk continuum and reflection component, we estimate a black hole mass of $10.5^{+7.7}_{-3.0}$, spin of $0.79^{+0.03}_{-0.13}$, and disk inclination angle of $42^{\circ}$$50^{\circ}$, which match well with the previously reported spectro-polarimetric measurements. The inferred source distance of $\sim 3.5$ kpc is consistent with the recent estimate based on optical spectroscopy. We find a weakly variable or stable disk and a highly variable Comptonized component.
@@
facilities: NICER, NuSTAR, AstroSat, Insight-HXMT, and MAXI.software: HEASoft (v.6.34; heasoft2014ascl.soft08004N), XSPEC (v.12.14.1; Arnaud1996ASPC..101...17A), CZTPIPELINE (v.3.0.1; http://astrosat-ssc.iucaa.in/cztiData), HXMTDAS (v.2.06; Zhao2020ASPC..527..469Z), Stingray (Huppenkothen2019JOSS....4.1393H; Huppenkothen2019ApJ...881...39H), Matplotlib (Hunter2007CSE.....9...90H).
1 Introduction
@@
Transient low-mass black hole X-ray binaries (BHXRBs) exhibit substantial evolution in their accretion geometry, particularly in the corona and accretion disk’s size and shape, across different spectral states during sporadic outbursts. These spectral state transitions are typically traced using the hardness–intensity diagram (HID), which follows a characteristic q-shaped track based on the changes in the source hardness with count rates in an anti-clockwise manner (Homan2001ApJS..132..377H; Belloni2005A&A...440..207B; HomanBelloni2005Ap&SS.300..107H; Fender2009MNRAS.396.1370F; Belloni2010LNP...794...53B; Belloni2011BASI...39..409B). During a typical outburst, BHXRBs spend most of their time in either the low/hard state (LHS) or the high/soft state (HSS), with relatively brief excursions through the short-lived HIMS and the SIMS.
@@
The accretion geometry in the HSS is now well understood, with a standard optically thick and geometrically thin accretion disk (Shakura1973A&A....24..337S) extending close to the innermost circular orbit (ISCO). This state is characterized by thermal disk emission peaking at $\approx 1$ keV and low levels of X-ray variability. In contrast, the LHS is dominated by Comptonized hard X-ray emission from a hot corona and exhibits strong variability. However, the geometry of the inner accretion flow in the LHS remains debated, particularly regarding whether the disk extends to the ISCO or is truncated at a large radius by a radiatively inefficient inner hot flow (2006ApJ...653..525M; 2008ApJ...679L.113M; 2006MNRAS.367..659D; 2008MNRAS.387.1489R; 2010MNRAS.407.2287D; 2014MNRAS.437..316K; Bambi2021SSRv..217...65B, for a review).
@@
The HIMS and SIMS are transitional phases, where significant contributions from both the disk and coronal emission are found in the observed spectrum. The HIMS typically shows relatively strong Comptonized emission and large variability, with often detection of type-C quasi-periodic oscillations (QPOs; Casella2004A&A...426..587C; Casella2005ApJ...629..403C). As the mass accretion rate increases, a source starts transitioning towards the SIMS, and the disk moves close to the ISCO (Done2007A&ARv..15....1D; Belloni2011BASI...39..409B). The observed spectrum is then dominated by the thermal emission from the disk, and the source variability becomes weaker in comparison to the HIMS. Type-B QPOs, which are relatively broader than the Type-C ones, are sometimes detected in the SIMS. Hence, the study of these intermediate states offers crucial insights into the dynamics of the inner accretion flow. However, BHXRBs do not always evolve through all four canonical spectral states during an outburst. In some cases, the source remains in the LHS throughout the entire outburst, likely due to a rapid change or insufficient rise in the mass accretion rate. Such events are referred to as hard-only or failed outbursts (Capitanio2009MNRAS.398.1194C; Stiele2016MNRAS.460.1946S; Sahu2024ApJ...975..165S).
@@
The hard X-ray emission observed in BHXRBs is commonly attributed to thermal Comptonization, which originates from the inverse Compton scattering of soft disk photons by hot electrons in the corona. The electron distribution is predominantly thermal or Maxwellian at a mildly relativistic temperature and shows a sharp cutoff at $\approx 100$ keV, particularly in the LHS (Done2007A&ARv..15....1D; Zdziarski2020MNRAS.492.5234Z). However, growing observational evidence for the presence of hard X-ray tails extending well beyond the thermal cutoff in several BHXRBs in the intermediate and soft spectral states suggests that the coronal electron population also includes a significant fraction of non-thermal electrons, with a power-law distribution (McConnell2000ApJ...543..928M; McConnell2002ApJ...572..984M; Wardzinski2002MNRAS.337..829W; Cadolle2006A&A...446..591C; Poutanen2009ApJ...690L..97P; Jourdain2012ApJ...744...64J; Cangemi2021A&A...650A..93C; Cangemi2023A&A...669A..65C). Similar non-thermal hard X-ray tail has also been reported in the LHS of the bright low-mass BHXRB MAXI J1820+070 (Zdziarski2021ApJ...914L...5Z). Such a hybrid electron distribution, comprising thermal and non-thermal electron populations, can significantly modify the shape of the Comptonized spectrum (Poutanen2009ApJ...690L..97P; Gierlinski1999MNRAS.309..496G; Zdziarski2001ApJ...554L..45Z). While the Comptonization due to thermally distributed electrons dominates the low-energy part of the spectrum, the high-energy excess beyond the thermal cutoff can be described by the Comptonization associated with the non-thermal electron distribution. Moreover, the non-thermal electron population can enhance electron-positron pair production, thereby introducing a thermostat effect that regulates and potentially lowers the coronal electron temperature (Zdziarski2021ApJ...914L...5Z).
@@
Another important component of the X-ray spectra of BHXRBs is the reflection emission, which arises due to the reflection of hard coronal X-rays by the disk. The reflection spectrum comprises mainly the iron $K_{\alpha}$ emission line, peaking at 676-7 keV, photoelectric absorption at low energies, and the reflection hump at 203020-30 keV originating due to Compton scattering of the high-energy photons. Gravitational redshift and the Doppler effect can distort the shape of the iron $K_{\alpha}$ line if the reflecting material is close to the ISCO (Fabian1989MNRAS.238..729F). The modeling of this reflection spectrum is one of the main approaches to estimate the spin of the black hole, independent of the black hole mass and source distance. Another widely used method for estimating black hole spin is the disk continuum modeling, which requires prior knowledge of the black hole mass and source distance. Both of these methods rely on the assumption that the accretion disk is at the ISCO or at least very close to it, which is usually observed in the soft state. Furthermore, the presence of a high-energy tail in the soft state, which produces the relativistic reflection features, allows us to constrain the black hole spin by modeling the reflection component and disk continuum together (Chand2022ApJ...933...69C; Zdziarski2024ApJ...962..101Z; Zdziarski2024ApJ...967L...9Z). A detailed overview of the different methods of spin estimation is provided in a recent review paper by Zdziarski2025arXiv250600623Z.
The bright low-mass X-ray transient Swift J1727.8–1613 was first detected on 2023 August 24 (Kennea2023GCN.34540....1K; Negoro2023GCN.34544....1N; Page2023GCN.34537....1P), with the source flux reaching a peak of $\sim 7$ Crab in the 155015–50 keV band (Palmer2023ATel16215....1P). Optical follow-up observations identified the source as a low-mass BHXRB residing at a distance of $3.4\pm 0.3$ kpc, with an orbital period of about $\sim 10$ hours (Mata2025A&A...693A.129M). Based on reflection modeling in the LHS, Peng2024ApJ...960L..17P reported a nearly maximally spinning black hole having a mass of $\sim 10M_{\odot}$ and a disk inclination of $\sim 40^{\circ}$. Later, Svoboda2024ApJ...966L..35S estimated the black hole spin to be $0.87\pm 0.03$ by jointly modeling the disk continuum and reflection component in the HSS. From polarimetric measurements, the authors suggested an inclination angle in the range of $30^{\circ}-50^{\circ}$. Using a new analytical model to fit the spatial structure of the jet, zdziarski2025ApJ...986L..35Z also favored a black hole mass below $13~M_{\odot}$ and a distance less than 5 kpc for this source. Additionally, a substantial variation in the polarization degree across different spectral states, implying an evolution in the inner accretion flow, is also observed in this source (Veledina2023ApJ...958L..16V; Svoboda2024ApJ...966L..35S). The polarization studies also suggest that the corona is radially extended in the disk plane and lies perpendicular to the jet axis (Veledina2023ApJ...958L..16V; Ingram2024ApJ...968...76I). Apart from this, Liu2024arXiv240603834L showed that the broadband spectrum of Swift J1727.8–1613 in the LHS is complex and requires Comptonization from a hybrid electron distribution. While the authors found the contribution from non-thermal electrons to be small in the LHS, it increased as the source moved toward the spectral transition. The QPO frequencies were also found to be evolving from $\sim 1.4$ to $\sim 2.6$ Hz over the HIMS (Nandi2024MNRAS.531.1149N) and $\sim 0.2$ Hz to $\sim 1.86$ Hz over the transitioning period from the intermediate to the soft states (Chatterjee2024ApJ...977..148C). Similar evolution in the QPO frequencies from $\sim 0.3$ Hz in the LHS to $\sim 7$ Hz in the HIMS, along with a shrinkage in the coronal shape and disk truncation radius from $30-40~\rm r_{g}$ to $10~\rm r_{g}$, was also reported using NICER observations (Rawat2025A&A...697A.229R).
@@
In this work, we carry out a detailed broadband spectral and timing study of the newly discovered low-mass BHXRB Swift J1727.8–1613 using multi-mission observations from NICER, NuSTAR, AstroSat, and Insight-HXMT during two distinct spectral states: the HIMS and SIMS. We demonstrate that the spectral data up to 78 keV in the HIMS requires models with two physically distinct thermal Comptonizing regions. Unlike previous studies in the LHS, our analysis of the extended broadband spectrum up to $200$ keV reveals that the hard X-ray tail in the HIMS becomes prominent only above $\sim 100$ keV. This spectral behavior cannot be accounted for by thermal Comptonization alone and instead requires a hybrid Comptonization model involving both thermal and non-thermal electron populations, underscoring the complex nature of the corona. We also show that the disk-corona geometry evolves significantly across the two observations. Utilizing broadband X-ray data in the SIMS, we estimate key parameters such as mass and spin of the black hole, disk inclination, and source distance using different models for the accretion disk, and examine the model dependence of these parameters. In addition, we investigate the variability properties of the source in both the HIMS and SIMS through a detailed analysis of the power density spectra (PDS) and rms–energy spectra. The paper is organized as follows. Section 2 describes the observations and data reduction. In Section 3, we present the results from our spectral and timing analysis. We discuss the implications of our findings in Section 4, and summarize our conclusions in Section 5.
2 Observations and Data Reduction
@@
We analyze two distinct sets of observations of the recently discovered low-mass BHXRB SWIFT J1727.8–1613, obtained during its 2023 outburst using NICER, NuSTAR, AstroSat, and Insight-HXMT. These data sets were likely taken during the intermediate states (see below) and remain unexplored in terms of broadband spectral modeling and detailed variability studies. The first dataset comprises simultaneous observations with NICER and NuSTAR, and is complemented by data from AstroSat and Insight-HXMT, providing broadband spectral coverage up to 200 keV near the peak of the outburst. The second dataset includes contemporaneous observations with NICER and NuSTAR, acquired during the decay phase of the outburst. Details of these observations are summarized in Table 1, and their positions within the outburst are indicated on the long-term MAXI light curve of the source, as shown in Figure 1.
@@
To obtain an initial insight into the spectral states of the source during the two observations, we derived HID using the MAXI observations of Swift J1727.8–1613 over the shaded interval in Figure 1, covering the entire 2023 outburst. We first calculated the hardness ratio (HR) by dividing the count rate in the 441010 keV band by that in the 242-4 keV band. The HID was then obtained by plotting the HR against the MAXI count rate in the 222020 keV band. Figure 2 displays the resulting HID, which indicates that the source underwent a full spectral transition during this outburst. The approximate positions of the two NICER observations analyzed in this study are overplotted on the HID based on their respective start times. These positions suggest that the source was likely in the HIMS during Obs1, characterized by an HR of $\sim 0.7$, and possibly transitioned to the SIMS during Obs2, with the HR decreasing to $\sim 0.3$.
2.1 NICER

### v2

@@ Abstract
We present a comprehensive broadband spectral and variability study of the newly detected black hole X-ray binary Swift J1727.8–1613 in the intermediate states during its 2023 outburst, using multi-mission observations from NICER, NuSTAR, AstroSat, and Insight-HXMT. Spectral data up to 78 keV in the hard-intermediate state (HIMS) require models with two Comptonizing regions. In contrast, models with a single Comptonizing region adequately describe the soft-intermediate states (SIMS), implying a significant evolution in the disk-corona geometry between the states. The hard X-ray tail above $100$ keV in the HIMS, detected with both AstroSat/CZTI and Insight-HXMT/HE, indicates that the electron population in the corona is not purely thermal but rather hybrid, with a power-law distribution above the thermal cutoff. While both the reflection modeling and disk continuum fitting favor a truncated disk geometry in the HIMS, the disk in the SIMS moves substantially closer to the innermost stable circular orbit, accompanied by a significant rise in disk temperature. This interpretation is further supported by the increase in the QPO frequency from $\sim 1.3$ to $\sim 6.6$ Hz. From joint modeling of the disk continuum and reflection component and assuming the distance of 3.4 kpc, we estimate a black hole mass of $10.3^{+5.5}_{-2.5}~M_{\odot}$, spin of $0.79^{+0.07}_{-0.15}$, and the disk inclination angle of $\sim 37\arcdeg$$53\arcdeg$, which match well with the previously reported spectro-polarimetric measurements. We find a weakly variable or stable disk and a highly variable Comptonized component.
@@
facilities: NICER, NuSTAR, AstroSat, Insight-HXMT, and MAXI.software: HEASoft (v.6.34; Nasa High Energy Astrophysics Science Archive Research Center (Heasarc), 2014), XSPEC (v.12.14.1; Arnaud, 1996), CZTPIPELINE (v.3.0.1; http://astrosat-ssc.iucaa.in/cztiData), HXMTDAS (v.2.06; Zhao, 2020), Stingray (Huppenkothen et al., 2019b, a), Matplotlib (Hunter, 2007).
I Introduction
@@
Transient low-mass black hole X-ray binaries (BHXRBs) exhibit substantial evolution in their accretion geometry, particularly in the corona and in the size and shape of the accretion disk, across different spectral states during sporadic outbursts. These spectral state transitions are typically traced using the hardness–intensity diagram (HID), which follows a characteristic q-shaped track based on the changes in the source hardness with count rates in an anti-clockwise manner (Homan et al., 2001; Belloni et al., 2005; Homan and Belloni, 2005; Fender et al., 2009; Belloni, 2010; Belloni et al., 2011). During a typical outburst, BHXRBs spend most of their time in either the low/hard state (LHS) or the high/soft state (HSS), with relatively brief excursions through the short-lived HIMS and the SIMS.
@@
The accretion geometry in the HSS is now well understood, with a standard optically thick and geometrically thin accretion disk (Shakura and Sunyaev, 1973) extending close to the innermost circular orbit (ISCO). This state is characterized by thermal disk emission peaking at $\approx 1$ keV and low levels of X-ray variability. In contrast, the LHS is dominated by Comptonized hard X-ray emission from a hot corona and exhibits substantial variability. However, the geometry of the inner accretion flow in the LHS remains debated, particularly regarding whether the disk extends to the ISCO or is truncated at a large radius by a radiatively inefficient inner hot flow (Miller et al., 2006, 2008; Done and Gierliński, 2006; Reis et al., 2008; Done and Diaz Trigo, 2010; Kolehmainen et al., 2014; Bambi et al., 2021, for a review).
@@
The HIMS and SIMS are transitional phases in which significant contributions from both disk and coronal emission are observed in the spectrum. The HIMS typically shows relatively strong Comptonized emission and strong variability, with a common detection of type-C quasi-periodic oscillations (QPOs; Casella et al., 2004, 2005). As the mass accretion rate increases, a source starts transitioning towards the SIMS, and the disk moves close to the ISCO (Done et al., 2007; Belloni et al., 2011). The observed spectrum is then dominated by thermal emission from the disk, and the source variability is weaker than in the HIMS. Type-B QPOs, which are relatively broader than Type-C QPOs, are sometimes detected in SIMS. Hence, the study of these intermediate states offers crucial insights into the dynamics of the inner accretion flow. However, BHXRBs do not always evolve through all four canonical spectral states during an outburst. In some cases, the source remains in the LHS throughout the entire outburst, likely due to a rapid change or insufficient rise in the mass accretion rate. Such events are referred to as hard-only or failed outbursts (Capitanio et al., 2009; Stiele and Yu, 2016; Sahu et al., 2024).
@@
The hard X-ray emission observed in BHXRBs is commonly attributed to thermal Comptonization, arising from the inverse-Compton scattering of soft-disk photons by hot electrons in the corona. The electron distribution is predominantly thermal or Maxwellian at a mildly relativistic temperature and shows a sharp cutoff at $\approx 100$ keV, particularly in the LHS (Done et al., 2007; Zdziarski et al., 2020). However, growing observational evidence for the presence of hard X-ray tails extending well beyond the thermal cutoff in several BHXRBs in the intermediate and soft spectral states suggests that the coronal electron population also includes a significant fraction of non-thermal electrons, with a power-law distribution (McConnell et al., 2000, 2002; Wardziński et al., 2002; Cadolle Bel et al., 2006; Poutanen and Vurm, 2009; Jourdain et al., 2012; Cangemi et al., 2021, 2023). Similar non-thermal hard X-ray tail has also been reported in the LHS of the bright low-mass BHXRB MAXI J1820+070 (Zdziarski et al., 2021c). Such a hybrid electron distribution, comprising thermal and non-thermal electron populations, can significantly modify the shape of the Comptonized spectrum (Poutanen and Vurm, 2009; Gierliński et al., 1999; Zdziarski et al., 2001). While Comptonization from a thermally distributed electron population dominates the low-energy spectrum, the high-energy excess beyond the thermal cutoff can be described by Comptonization associated with the non-thermal electron distribution. Moreover, the non-thermal electron population can enhance electron-positron pair production, thereby introducing a thermostat effect that regulates and potentially lowers the coronal electron temperature (Zdziarski et al., 2021c).
@@
Another critical component of the X-ray spectra of BHXRBs is the reflection emission, which arises due to the reflection of hard coronal X-rays by the disk. The reflection spectrum comprises mainly the iron $K_{\alpha}$ emission line, peaking at 6–7 keV, photoelectric absorption at low energies, and the reflection hump at 20–30 keV, originating from Compton scattering of high-energy photons. Gravitational redshift and the Doppler effect can distort the shape of the iron $K_{\alpha}$ line if the reflecting material is close to the ISCO (Fabian et al., 1989). Modeling this reflection spectrum is a primary approach for estimating the black hole spin, independent of mass and source distance. Another widely used method for estimating black hole spin is the disk continuum modeling, which requires prior knowledge of the black hole mass and source distance. Both methods rely on the assumption that the accretion disk is at the ISCO or very close to it, a configuration typically observed in the soft state. Furthermore, the presence of a high-energy non-thermal tail in the soft state irradiates the inner accretion disk. It produces relativistic reflection features, including a broadened Fe $K_{\alpha}$ line, whose degree of broadening depends on the black hole spin. Joint modeling of the relativistic reflection features and the thermal disk continuum, whose inner radius is linked to the ISCO, helps break degeneracies between different parameters, such as black hole mass, distance, and disk inclination angle, thereby improving constraints on the black hole spin (Parker et al., 2016; Chand et al., 2022; Zdziarski et al., 2024a, b). A detailed overview of the different methods of spin estimation is provided in a recent review paper by Zdziarski et al. (2026).
The bright low-mass X-ray transient Swift J1727.8–1613 was first detected on 2023 August 24 (Kennea and Swift Team, 2023; Negoro et al., 2023; Page et al., 2023), with the source flux reaching a peak of $\sim 7$ Crab in the 155015–50 keV band (Palmer and Parsotan, 2023). Optical follow-up observations identified the source as a low-mass BHXRB with an orbital period of about $\sim 10$ hours (Mata Sánchez et al., 2025). The distance was estimated to be $D=3.4\pm 0.3$ kpc of Mata Sánchez et al. (2025) and $5.5^{+1.4}_{-1.1}$ kpc by Burridge et al. (2025). On the other hand, Zdziarski et al. (2025) argued for the distance $\lesssim$5 kpc by considering the Eddington limit on the bolometric hard-state luminosity. Here, we adopt the distance of Mata Sánchez et al. (2025), but consider the scalings of our derived parameters with $D$.
Based on reflection modeling in the LHS, Peng et al. (2024) reported a nearly maximally spinning black hole having a mass of $\sim 10M_{\odot}$ and a disk inclination of $\sim 40\arcdeg$. Later, Svoboda et al. (2024) estimated the black hole spin to be $0.87\pm 0.03$ by jointly modeling the disk continuum and reflection component in the HSS. Based on polarimetric measurements, the authors suggested an inclination angle of $30\arcdeg$$50\arcdeg$. Additionally, substantial variation in the degree of polarization across spectral states, implying evolution in the inner accretion flow, is observed in this source (Veledina et al., 2023; Svoboda et al., 2024). The polarization studies also suggest that the corona is radially extended in the disk plane and lies perpendicular to the jet axis (Veledina et al., 2023; Ingram et al., 2024).
[…]
@@
In this work, we carry out a detailed broadband spectral and timing study of the newly discovered low-mass BHXRB Swift J1727.8–1613 using multi-mission observations from NICER, NuSTAR, AstroSat, and Insight-HXMT during two distinct spectral states: the HIMS and SIMS. We demonstrate that the spectral data up to 78 keV in the HIMS requires models with two physically distinct thermal Comptonizing regions. Unlike previous studies in the LHS, our analysis of the extended broadband spectrum up to $200$ keV reveals that the hard X-ray tail in the HIMS becomes prominent only above $\sim 100$ keV. This spectral behavior cannot be accounted for by thermal Comptonization alone and instead requires a hybrid Comptonization model that incorporates both thermal and non-thermal electron populations, underscoring the corona’s complexity. We also show that the disk-corona geometry evolves significantly across the two observations. Using broadband X-ray data in SIMS, we estimate key parameters such as the black hole’s mass and spin, and the disk inclination angle, using different accretion disk models and examine the model dependence of these parameters. In addition, we investigate the variability properties of the source in both HIMS and SIMS through detailed analyses of power density spectra (PDS) and rms–energy spectra. The paper is organized as follows. Section II describes the observations and data reduction. In Section III, we present the results from our spectral and timing analysis. We discuss the implications of our findings in Section IV, and summarize our conclusions in Section V.
II Observations and Data Reduction
@@
We analyze two distinct sets of observations of the recently discovered low-mass BHXRB SWIFT J1727.8–1613, obtained during its 2023 outburst using NICER, NuSTAR, AstroSat, and Insight-HXMT. These datasets were obtained during the intermediate states (see below) and remain unexplored for broadband spectral modeling and detailed variability studies. The first dataset comprises simultaneous observations with NICER and NuSTAR. It is complemented by data from AstroSat and Insight-HXMT, providing broadband spectral coverage up to 200 keV near the peak of the outburst. The second dataset includes contemporaneous observations with NICER and NuSTAR, acquired during the decay phase of the outburst. Details of these observations are summarized in Table 1, and their positions within the outburst are indicated on the long-term MAXI light curve of the source, as shown in Figure 1.
@@
To obtain an initial insight into the spectral states of the source during the two observations, we derived HID using the MAXI observations of Swift J1727.8–1613 over the shaded interval in Figure 1, covering the entire 2023 outburst. We first calculated the hardness ratio (HR) by dividing the count rate in the 441010 keV band by that in the 242-4 keV band. The HID was then obtained by plotting the HR against the MAXI count rate in the 222020 keV band. Figure 2 displays the resulting HID, which indicates that the source underwent a complete spectral transition during this outburst. The approximate positions of the two NICER observations analyzed in this study are overplotted on the HID based on their respective start times. These positions suggest that the source was likely in the HIMS during Obs1, characterized by an HR of $\sim 0.7$, and possibly transitioned to the SIMS during Obs2, with the HR decreasing to $\sim 0.3$.
II.1 NICER